Solve the following system of ODE

(40D+4)y - (2D+1)z = e^-x

(D+8)y - 3z = 5e^-x ; D=d/dx


1
Expert's answer
2021-02-03T01:24:40-0500

Solve the following system of ODE

(40D+4)y - (2D+1)z = e^-x

(D+8)y - 3z = 5e^-x ; D=d/dx

"\begin{cases}
 40\frac{dy}{dx}+4y-2\frac{dz}{dx}-z=e^{-x} \\
 \frac{dy}{dx}+8y-3z=5e^{-x} 
\end{cases}"

Solution:

From second equation:

"z=\frac13(\frac{dy}{dx}+8y-5e^{-x} )" (*)

"\frac{dz}{dx}=\frac13(\frac{d^2y}{dx^2}+8\frac{dy}{dx}+5e^{-x})"

Substitute "z" and "\frac{dz}{dx}" into the first equation:

"40\frac{dy}{dx}+4y-\frac23(\frac{d^2y}{dx^2}+8\frac{dy}{dx}+5e^{-x})-\frac13(\frac{dy}{dx}+8y-5e^{-x} )=e^{-x}"

"-\frac23\frac{d^2y}{dx^2}+\frac{103}{3}\frac{dy}{dx}+\frac43y=\frac83e^{-x}"

"2\frac{d^2y}{dx^2}-103\frac{dy}{dx}-4y=-8e^{-x}" (**)

First find the general solution of the corresponding homogeneous equation:

"2\frac{d^2y_0}{dx^2}-103\frac{dy_0}{dx}-4y_0=0"

Let's compose and solve the characteristic equation:

"2\lambda^2-103\lambda-4=0"

"\lambda_1=\frac{103-\sqrt{10641}}{4}" , "\lambda_2=\frac{103+\sqrt{10641}}{4}" .

"y_0=C_1e^{\lambda_1x}+C_2e^{\lambda_2x}=C_1e^{\frac{103-\sqrt{10641}}{4}x}+C_2e^{\frac{103+\sqrt{10641}}{4}x}" .

Find a typical (specific) solution of the non-homogeneous equation in the form "y_p=Ae^{-x}"

Derivatives of the solution: "\frac{dy_p}{dx}=-Ae^{-x}" , "\frac{d^2y_p}{dx^2}=Ae^{-x}"

Substitute those "solutions" into the (**):

"2Ae^{-x}+103Ae^{-x}-4Ae^{-x}=-8e^{-x}"

"2A+103A-4A=-8"

"A=-\frac{101}{8}"

"y_p=-\frac{101}{8}e^{-x}" .

Add the typical and the complementary solutions to get the complete solution:

"y=y_0+y_p=C_1e^{\frac{103-\sqrt{10641}}{4}x}+C_2e^{\frac{103+\sqrt{10641}}{4}x}-\frac{101}{8}e^{-x}" .

To find "z" substitute "y" into the (*):

"z=\frac13(C_1\frac{103-\sqrt{10641}}{4}e^{\frac{103-\sqrt{10641}}{4}x}+C_2\frac{103+\sqrt{10641}}{4}e^{\frac{103+\sqrt{10641}}{4}x}+\frac{101}{8}e^{-x}+"

"8(C_1e^{\frac{103-\sqrt{10641}}{4}x}+C_2e^{\frac{103+\sqrt{10641}}{4}x}-\frac{101}{8}e^{-x})-5e^{-x} )="

"C_1\frac{135-\sqrt{10641}}{12}e^{\frac{103-\sqrt{10641}}{4}x}+C_2\frac{135+\sqrt{10641}}{12}e^{\frac{103+\sqrt{10641}}{4}x}-\frac{249}{8}e^{-x}"

Answer: "\begin{cases}
 y=C_1e^{\frac{103-\sqrt{10641}}{4}x}+C_2e^{\frac{103+\sqrt{10641}}{4}x}-\frac{101}{8}e^{-x} \\
 z=C_1\frac{135-\sqrt{10641}}{12}e^{\frac{103-\sqrt{10641}}{4}x}+C_2\frac{135+\sqrt{10641}}{12}e^{\frac{103+\sqrt{10641}}{4}x}-\frac{249}{8}e^{-x}
\end{cases}"